OpenStudy (bg256):

in Δ ABC, AB=x, BC=y, and CA=2x. A similarity transformation with a scale factor of 0.5 maps Δ ABC to Δ MNO, such that vertices M, N, and O correspond to A, B, and C, respectively. If OM=5, what is AB?

OpenStudy (bg256):

A. AB=2.5 B. AB= 10 C. AB= 5 D. AB= 1.25 E. AB= 2

OpenStudy (welshfella):

corresponding sides would be in the ratio 1:0.5 withe MNO being smallest triangle. which side in triangle ABC corresponds to OM in triangle MNO?

OpenStudy (bg256):

Side CA I believe @welshfella

OpenStudy (welshfella):

thats correct

OpenStudy (welshfella):

so OM = 0.5*CA and CA = 2x so OM = 0.5* 2x also OM = 5 so 5 = 0.5* 2x can you continue?

OpenStudy (welshfella):

do you follow that OK?

OpenStudy (welshfella):

each of the sides in MNO are 1/2 length of corresponding sides in ABC - because the scale factor is 0.5.

OpenStudy (welshfella):

Now as AB = x if you solve the equation for x that's the length of AB.

OpenStudy (bg256):

I believe so I had to think a bit after that but if I did it right wouldn't AB =5? @welshfella

OpenStudy (welshfella):

yes x = 5 so AB = 5

OpenStudy (bg256):

Okay thanks for the help @welshfella